Find the mass and center of mass of a triangular lamina with vertices
(0, 0), (1, 0)
and
(0, 2)
if the density function is p(x,y)=1+x+y
1
Expert's answer
2020-08-13T18:36:22-0400

Equation of the hypotenuse of the right angle triangle is "y = -2x + 2"

Mass of the plate is given by

"M = \int_0^1 \int_0^{(2-2x)} (1+x+y)dydx"


"\int_0^1 (y+xy+\frac{y^2}{2})|_0^{2-2x} dx = \int_0^1 4-4x dx = 2"



Center of mass,

X-coordinate,

"M_x = \frac{1}{M} \int_0^1 \int_0^{2-2x} x(1+y+x)dydx = \frac{1}{M} \int_0^1 [xy+x^2y+\frac{xy^2}{2} ]_0^{2-2x} dx"


"= \frac{1}{M} \int_0^1 (4x-4x^2) dx = \frac{1}{3}"


Y-coordinate,

"M_y=\frac{1}{M} \int_0^1 \int_0^{2-2x} y(1+y+x)dydx"


"= \frac{1}{M} \int_0^1[ \frac{(2-2x)^2}{2}+\frac{x(2-2x)^2}{2} + \frac{(2-2x)^3}{3} ]dx"


"= \frac{3}{4}"